OpenStudy (anonymous):

Which is equal to 4^2 • 4^8? A. 4^16 B. 4^10 C. 4^6 D. 4^−6

OpenStudy (anonymous):

A?

OpenStudy (phebe):

what do you think?

Nnesha (nnesha):

when we multiply same bases we should `add` their exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

OpenStudy (phebe):

First thing you should do is solve both

OpenStudy (anonymous):

i think it is b.

Nnesha (nnesha):

right 8+2 =10

OpenStudy (anonymous):

your right

OpenStudy (phebe):

B is the correct answer

OpenStudy (phebe):

^^

OpenStudy (anonymous):

thanks everyone :)

OpenStudy (phebe):

Welcome

Nnesha (nnesha):

yw :=)